### v2

# Best-Worst Disaggregation: An approach to the preference disaggregation problem

Abstract
Preference disaggregation methods in Multi-Criteria Decision Analysis (MCDA) often encounter challenges related to inconsistency and cognitive biases when deriving a value function from experts’ holistic preferences. This paper introduces the Best-Worst Disaggregation (BWD) method, a novel approach that integrates the principles of the Best-Worst Method (BWM) into the disaggregation framework to enhance the consistency and reliability of derived preference models. BWD employs the “consider-the-opposite” strategy from BWM, allowing experts to provide two opposite pairwise comparison vectors of alternatives. This approach reduces cognitive load and mitigates anchoring bias, possibly leading to more reliable criteria weights and attribute value functions. An optimization model is formulated to determine the most suitable additive value function to the preferences expressed by an expert. The method also incorporates a consistency analysis to quantify and improve the reliability of the judgments. Additionally, BWD is extended to handle interval-valued preferences, enhancing its applicability in situations with uncertainty or imprecise information. We also developed an approach to identify a reference set, which is used for pairwise comparisons to elicit the value functions and weights. A case study in logistics performance evaluation demonstrates the practicality and effectiveness of BWD, showing that it produces reliable rankings aligned closely with experts’ preferences.
Keywords: Multi-Criteria Decision Analysis (MCDA); Best-Worst Method (BWM); Preference disaggregation; Consistency analysis; Imprecise preference information
1 Introduction
The field of Multi-Criteria Decision Analysis (MCDA) encompasses a wide range of methods designed to aid experts in evaluating a set of alternatives characterized by a set of attributes/criteria (Greco et al., , 2016). Within MCDA, value-based methods, such as Multi-Attribute Value Theory (MAVT) (Keeney and Raiffa, , 1993), Analytic Hierarchy Process (AHP) (Saaty, , 1977), and Best-Worst Method (BWM) (Rezaei, , 2015) , aim to assign a numerical value to each alternative. In these methods, experts assess the relative importance of criteria and the performance of alternatives on these criteria, often using pairwise comparisons to derive weights and scores. The scores are then aggregated to produce an overall ranking, emphasizing the total utility or value derived from each alternative.
Value-based methods are further divided into aggregation and disaggregation approaches. Aggregation methods, such as MAVT, elicit value functions and trade-offs among the objectives to evaluate and rank the alternatives. Disaggregation methods, including the UTA family (Siskos et al., , 2016), on the other hand, are appropriate when an expert can provide holistic preferential information on a reference set of alternatives. These methods will then help infer the underlying value functions and trade-offs. Doumpos and Zopounidis, (2011) provided a comprehensive review of preference disaggregation methods, emphasizing their connections with statistical learning techniques, and highlighting how these approaches can be integrated to enhance decision support systems. In this context, preference learning emerges as a promising alternative, functioning as an explainable AI that learns and maintains fully interpretable preferences. Preference disaggregation techniques focus on deriving preference structures that are both accurate and transparent, aligning with the growing emphasis on explainable models in decision-making (Rudin, , 2019). By ensuring that the learned preferences are easily understandable, preference learning enhances the usability and acceptance of MCDA methods.
Despite these methodological advances, a fundamental challenge remains. Behavioral decision theory suggests that decision-makers often exhibit inconsistencies in their judgment and are influenced by various cognitive biases (Tversky and Kahneman, , 1974). While disaggregation methods are powerful and widely applicable, their reliance on a single round of holistic comparisons may leave room for judgment inconsistencies to influence the results. The comparison processes might introduce some variability in ranking outcomes, especially when experts’ judgments are internally inconsistent, potentially affecting the stability of the inferred value functions (Montibeller and Von Winterfeldt, , 2015; Morton and Fasolo, , 2009). This consideration is particularly important for methods like the UTA family, most of which inherently assume a high level of consistency and rationality in the provided preference rankings (Vetschera, , 2006; Schilling et al., , 2007; Korhonen et al., , 2012). GRIP (Generalized Regression with Intensities of Preferences) (Figueira et al., , 2009) enriches the UTA family by explicitly modeling preference intensities through a regression framework, thereby offering analysts a more nuanced view of stakeholder judgments; this additional expressive power naturally comes with a need for some statistical expertise and appropriate software support.Barbati et al., (2024) proposed a Deck-of-Cards-based Robust Ordinal Regression (DOR) to push further the use of cardinal preferential information in preference disaggregation methods.
One practical issue that can arise when applying UTA-type methods is that incomplete preference information may leave the analyst with several—occasionally infinitely many—additive value models that all fit the decision maker’s holistic judgments (Greco et al., , 2008). This plurality, while faithful to the data, may translate into more than one plausible recommendation. Eliciting additional judgments is an effective way to narrow the solution set, but time or cognitive-load constraints sometimes limit how much extra information can be gathered (Ciomek et al., , 2017; Salo and Hämäläinen, , 2010). Robust Ordinal Regression and related robustness-analysis procedures offer valuable insight into the full spectrum of admissible rankings, yet presenting a range of models instead of a single solution can feel abstract to users who prefer an immediately actionable outcome. Regularization frameworks—designed to balance goodness-of-fit with smoothness of the inferred value functions—provide another principled way to identify a representative model (Doumpos and Zopounidis, , 2007; Liu et al., , 2019), although their implementation typically demands a higher level of modeling expertise (Kadziński et al., , 2017).
To further address these challenges, this study introduces a novel solution called Best-Worst Disaggregation (BWD) method, which integrates the idea of BWM within the disaggregation framework of MCDA. Let us note that the idea of combining the two approaches was already suggested by Barbati et al., (2024). The BWD method leverages the structured pairwise comparison approach of BWM alongside the foundational principles of disaggregation methods to derive the value functions, reliably and consistently, from the preferences provided by the experts. The proposed method adopts the consider-the-opposite strategy (Mussweiler and Strack, , 2000; Mussweiler et al., , 2000), by involving comparisons between two opposite reference points—best and worst—and other criteria. This strategy has been incorporated into numerous psychological studies to enhance judgment consistency and demonstrated its efficacy to mitigate the negative effects of the anchoring bias (Adame, , 2016; Joslyn et al., , 2011; Mussweiler, , 2002; Rezaei, , 2021; Rezaei et al., , 2022, 2024). While DOR is expressly designed to capture both ranking and preference intensity—thanks to its combination of the Deck-of-Cards visual metaphor and ordinal-regression machinery—this richer elicitation protocol is best suited to decision settings where stakeholders have the time and inclination for in-depth interaction.
We summarize the contribution of this study as follows: Introducing a preference disaggregation method by combining the Best-Worst Method (BWM) with the disaggregation framework in MCDA, employing the “consider-the-opposite” strategy to reduce cognitive bias, and developing an optimization model to derive the most suitable additive value function based on experts’ pairwise comparisons. Extending the method to handle interval-valued preferences broadens its applicability in situations with uncertainty or imprecise information. A systematic approach is proposed for selecting a reference set that covers all criteria levels, minimizes cognitive load, and avoids Pareto dominance. Consistency analysis is incorporated using both ordinal and cardinal consistency ratios to ensure higher reliability of results, along with sensitivity analysis to assess the maximum and minimum possible ranks for alternatives, evaluating their stability.
The remainder of this paper is structured as follows: Section 2 provides preliminary information on the underlying theories of MCDA, focusing on BWM and UTA methods. Section 3 details the proposed BWD method, illustrating how it integrates the structured pairwise comparison approach of BWM with the disaggregation framework of UTA to derive value functions. In its subsections we explore consistency and sensitivity analysis as well as interval-valued extensions. Section 4 discusses a real-world application of the method, showcasing its practical utility. Finally, Section 5 concludes the paper and suggests directions for future research.
2 Preliminaries
2.1 Multi-attribute value theory (MAVT)
Multi-Attribute Value Theory (MAVT) is a framework in MCDA that assumes that each alternative can be characterized by a list of its relevant attributes. A typical MCDA problem consists of a non-empty finite set of $m$ alternatives whose indices range in the set $M=\{1,\ldots,m\}$ and a finite set of $n$ attributes (also known as criteria) indexed by the set $N=\{1,\ldots,n\}$. The set of possible levels that can be achieved by a generic alternative with respect to the $j$th attribute is denoted by $X_{j}$. Assuming all relevant attributes have been considered, each alternative can be associated with a consequence vector $\mathbf{x}=(x_{1},\ldots,x_{n})\in X_{1}\times\cdots\times X_{n}$, where $x_{j}\in X_{j}$ represents the level of the $j$th attribute achieved by the alternative. For simplicity, we will consider consequence vectors $\mathbf{x}$ in place of alternatives. In MAVT, each attribute can be rescaled and normalized into the interval $[0,1]$ using a function $v_{j}:X_{j}\rightarrow[0,1]$ where we assign values 0 and 1 to the least $\underline{x}_{j}$ and most $\overline{x}_{j}$ desirable attribute levels, respectively. Hence, $v_{j}(x_{j})\in[0,1]$ can be interpreted as the value of the consequence evaluated with respect to the $j$th attribute. The primary challenge in MAVT is to find a function $V:[0,1]^{n}\rightarrow[0,1]$ that can correctly aggregate the $n$ values $v_{j}(x_{j})$ and represent the preferences of a decision-maker or an expert. Namely, given a set of alternatives, each of which can be uniquely represented by an associated consequence vector, the value function $V$ should be coherent with the preorder structure on the set of alternatives:
$V(\mathbf{x})\geq V(\mathbf{y})\Leftrightarrow\mathbf{x}\succsim\mathbf{y}\;\;\text{($\mathbf{x}$ is weakly preferred to $\mathbf{y}$)}$ (1)
Several theoretical results from the literature help restrict the set of compatible value functions. Under mild conditions such as preference independence and difference independence (Smith and Dyer, , 2021), there exist attribute value functions $v_{j}:X_{j}\rightarrow[0,1]$ such that:
$V(\mathbf{x})=\sum_{j=1}^{n}w_{j}v_{j}(x_{j})$ (2)
where the scaling constants (also called weights) are such that $w_{j}\geq 0$ for all $j$, and $w_{1}+\cdots+w_{n}=1$. This type of value function is called additive value function. Assuming this formulation, the main task of multi-attribute value theory methods is to find attribute value functions $v_{j}$ and scaling constants $w_{j}$. This is where we can roughly divide MAVT methods into two categories.
[…]
The second family of methods uses elicitation techniques on experts to find compatible attribute value functions $v_{j}$’s and scaling constants $w_{j}$’s. That is, these methods work entirely on the left-hand side of equivalence (1). Here we have combinations of methods like the mid-value splitting technique and the trade-off method (Eisenführ et al., , 2010), the SMART (Edwards and Barron, , 1994) and Swing families (Von Winterfeldt and Edwards, , 1986) as well as, among others, the already mentioned AHP and BWM.
2.2 Best-Worst Method
The Best-Worst Method (BWM) is a pairwise comparison-based method used to elicit the weights of the criteria in a structured way. The pairwise comparison data is collected from the expert who is asked to first identify a set of relevant decision criteria. Then the expert should identify the ‘best’ and the ‘worst’ criteria. The ‘best’ criterion is the one which, according to the expert, has the largest contribution to the goal of the decision-making problem, while the ‘worst’ has the smallest contribution. Then, the expert compares the best criterion, $B$, to all the others, and also all the criteria to the worst one $W$, using a number between 1 to 9 (or possibly some other scales). Finally, an optimization model is formulated and solved to find the weights, $w_{j}$, such that the maximum absolute difference between the pairwise comparisons and their associated weight ratios is minimized. The optimization model is as follows (Rezaei, , 2015),
$\underset{\xi,w_{1},\ldots,w_{n}}{\text{minimize}}$ $\xi$ (3)
subject to $\left|\frac{w_{B}}{w_{j}}-a_{Bj}\right|\leq\xi\hskip 14.22636pt\forall j\in N\setminus\{B\}$
$\left|\frac{w_{j}}{w_{W}}-a_{jW}\right|\leq\xi\hskip 14.22636pt\forall j\in N\setminus\{W\}$
$w_{1}+\cdots+w_{n}=1$
$w_{j}\geq 0\hskip 14.22636pt\forall j\in N\,.$
where $a_{Bj}$ represents the value of the comparison between the best criterion $B$ and the $j$th criterion, and similarly $a_{jW}$ is the comparison between the $j$th criterion and the worst one, $W$. Optimization problem (3) is non-linear and the optimal value of the objective function can be interpreted as a quantification of inconsistency. There exists a linear version of the BWM which is easier to solve, yet with very similar final results, which is presented as follows (Rezaei, , 2016).
$\underset{\xi^{L},w_{1},\ldots,w_{n}}{\text{minimize}}$ $\xi^{L}$ (4)
subject to $|w_{B}-a_{Bj}w_{j}|\leq\xi^{L}\hskip 14.22636pt\forall j\in N\setminus\{B\}$
$|w_{j}-a_{jW}w_{W}|\leq\xi^{L}\hskip 14.22636pt\forall j\in N\setminus\{W\}$
$w_{1}+\cdots+w_{n}=1$
$w_{j}\geq 0\hskip 14.22636pt\forall j\in N\,.$
Solving (4) results in the optimal weights and the optimal objective value can again be used as an inconsistency index. A survey on extensions and applications of the BWM method was proposed by Mi et al., (2019). In the BWM, questions on pairs on criteria regard their relative importance in the decision process. Conversely, in MAVT, ratios between weights are interpreted as rates of substitutions. To fill this gap, the Best-Worst Tradeoff (BWT) was introduced (Liang et al., , 2022) to adapt the BWM to the context of MAVT. Whereas the interpretation between the values of the comparisons used in BWM and BWT are different, the formulations of the models are sufficiently similar so that the former is still representative of the latter.
2.3 Preference disaggregation analysis framework
Preference disaggregation is an approach within MCDA that focuses on deriving criteria weights and value functions from some holistic preferences of experts on a reference set of alternatives (Doumpos et al., , 2022). This approach is particularly valuable when:
  • experts or decision-makers find it more practical to provide overall evaluations of alternatives rather than detailed information on criteria. An example is the application of preference learning to the selection of validators in a blockchain environment (Gehrlein et al., , 2023).
  • there is a sufficiently large set of well-known alternatives from which an expert or a decision-maker can easily extract a reference set on which she feels sufficiently confident to express her preferences. An example is the study by Corrente et al., (2013) on the levels of innovation of European countries.
  • comparing some pairs of criteria is difficult. For example, when it comes to choosing a holiday destination, we may be able to easily rank alternatives and yet we may not be able to assess the substitution rate between factors such as distance from home and time taken for the planning.
The UTA (UTilités Additives) method, proposed by Jacquet-Lagreze and Siskos, (1982), is a regression-based approach within the field of MCDA. This method was designed to infer additive value functions that align with the preferences of an expert, as expressed through the ranking of a set of alternatives. The core idea is to disaggregate these preferences to construct value functions that best replicate the given ranking, ensuring consistency across the set of alternatives. The goal of the UTASTAR procedure–an improved variant of the basic UTA method–is to derive additive value functions that reflect the expert’s holistic preferences. First, the expert provides a ranking on a reference set. Next, a linear optimization problem is solved to derive the attribute value functions for each criterion, ensuring that the attribute value functions replicate the provided rankings as closely as possible. More formally, the steps of UTASTAR are the following.
Concept and framework
In a typical multi-criteria decision-making problem, a set of $m$ alternatives $\{\mathbf{x}_{1}\ldots,\mathbf{x}_{m}\}$ has to be evaluated with respect to $n$ criteria. The UTA method seeks to derive a global value function. To simplify the search for such a function, and knowing that $V$ is unique up to an affine transformation, UTASTAR considers the weights implicitly in the global value function. That is, attribute value functions are normalized in such a way that
$\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1.$ (5)
By doing so, the explicit mention to weights can be omitted, as they simply correspond to the values $v_{j}(\overline{x}_{j})$ and the value of an alternative $\mathbf{x}$ becomes $V(\mathbf{x})=\sum_{j=1}^{n}v_{j}(x_{j})$.
Linear Programming and Inference
Jacquet-Lagreze and Siskos, (1982) proposed to use piecewise linear approximations of the attribute value functions $v_{j}$. For the $j$th criterion, the interval $[\underline{x}_{j},\overline{x}_{j}]$ can be divided by $(s_{j}-1)$ (with $s_{j}\in\mathbb{N}_{\geq 1}$) equally spaced points in $s_{j}$ subintervals, and the approximate attribute value function $v_{j}^{L}$ can be defined as follows. If $x_{j}\in[x_{j}^{k},x_{j}^{k+1}]$, then
$v_{j}^{L}(x_{j})=v_{j}^{L}(x_{j}^{k})+\frac{x_{j}-x_{j}^{k}}{x_{j}^{k+1}-x_{j}^{k}}\left[v_{j}^{L}(x_{j}^{k+1})-v_{j}^{L}(x_{j}^{k})\right]$ (6)
Very often one can also assume monotonicity, so that each piecewise linear attribute value function $v_{j}^{L}(x_{j})$ satisfies the following constraints (assuming that $j$ is a benefit criterion):
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{(k+1)})\;\;\forall j\text{ and }\forall k.$ (7)
The value function resulting from the aggregation of the piecewise linear attribute value functions is, then,
$V_{L}(\mathbf{x})=\sum_{j=1}^{n}v_{j}^{L}(x_{j})$ (8)
[…]
$\begin{cases}V_{L}(\mathbf{x})-V_{L}(\mathbf{y})\geq\delta&\text{if }\mathbf{x}\succ\mathbf{y}\\ V_{L}(\mathbf{x})-V_{L}(\mathbf{y})=0&\text{if }\mathbf{x}\sim\mathbf{y}\end{cases}$ (9)
[…]
$\tilde{V}(\mathbf{x})=\sum_{j=1}^{n}v_{j}^{L}(x_{j})-\sigma_{\mathbf{x}}^{+}+\sigma_{\mathbf{x}}^{-}$ (10)
where $x_{j}$ is the attribute level achieved by alternative $\mathbf{x}$ with respect to the $j$th attribute and $\sigma_{\mathbf{x}}^{+}$ and $\sigma_{\mathbf{x}}^{-}$ are overestimation and underestimation errors, respectively. Albeit not a strict requirement, pragmatism suggests that the cardinality of $R$ be significantly smaller than the cardinality of $M$.
Optimization
[…]
$\theta=$ minimize $\sum_{i\in R}\left(\sigma_{\mathbf{x}_{i}}^{+}+\sigma_{\mathbf{x}_{i}}^{-}\right)$ (11)
subject to $\tilde{V}(\mathbf{x}_{p})-\tilde{V}(\mathbf{x}_{q})\geq\delta\text{ if }\mathbf{x}_{p}\succ\mathbf{x}_{q}$
$\tilde{V}(\mathbf{x}_{p})-\tilde{V}(\mathbf{x}_{q})=0\text{ if }\mathbf{x}_{p}\sim\mathbf{x}_{q}$
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\text{ and }\forall k$
$\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
$v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N$
$\sigma_{\mathbf{x}_{i}}^{+},\sigma_{\mathbf{x}_{i}}^{-}\geq 0\hskip 14.22636pt\forall i\in R$
If the optimal value of the objective function is $\theta=0$, then the constructed attribute value functions can reflect the expert’s preferences. Conversely, if $\theta>0$, then the attribute value functions fail to reflect them and thus are only approximations.
Stability
Given the possible existence of multiple optimal or near-optimal solutions, it is suggested to take the average of the additive value functions that maximize the objective functions $v_{j}(\overline{x}_{j})$ for all $j\in N$, with the same constraints as in (11) and the addition of
$\sum_{i\in R}\left(\sigma_{\mathbf{x}_{i}}^{+}+\sigma_{\mathbf{x}_{i}}^{-}\right)\leq\theta^{*}+\varepsilon$
where $\theta^{*}$ is the optimum of (11) and $\varepsilon>0$ is extremely small, so that only the set of optimal or near-optimal solutions is considered.
Over the years, several variants and extensions of UTA methods have been developed to enhance its applicability and robustness.For instance, the UTADIS variant caters to sorting problems (Zopounidis and Paraschou, , 2013). Another notable extension is the Quasi-UTA model, which simplifies the process by using recursive exponential functions to reduce the information needed to build the utility function (Beuthe et al., , 2000). Since its inception, UTA methods have found applications across various fields (Dias et al., , 2018), including commercial strategy, venture capital investment evaluation, country risk assessment, business financing, and portfolio management (Siskos et al., , 2016). Methods such as UTAGMS and GRIP incorporate robust ordinal regression to handle interacting criteria and enhance the decision-making process (Greco et al., , 2008). These advancements ensure that the UTA methods remain relevant and powerful tools in the field of MCDA (Doumpos et al., , 2022).
3 Best-worst disaggregation method
Suppose we have a set of $m$ alternatives, represented by the profile vectors $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}$, which are evaluated with respect to a set of $n$ criteria. The performance matrix can be represented as follows,
$\mathbf{X}=\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{m}\end{pmatrix}=\begin{pmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\ddots&\vdots\\ x_{m1}&\cdots&x_{mn}\end{pmatrix}$
where the $i$th row is the profile of the $i$th alternative, i.e., $\mathbf{x}_{i}$. Following the precepts of UTASTAR, we identify a reference set of alternatives and we call $R$ its index set. Then, according to the principles of BWM, we identify the indices of the best and the worst alternatives in the reference set and we label them with the letters $B$ and $W$, respectively. Let us incidentally note that the freedom in the definition of the reference set $R$ allows the decision maker to choose it in such a way to simplify the choice of the best and the worst alternative within it. We then conduct a holistic comparison between the best alternative $\mathbf{x}_{B}$ and the other alternatives, as well as between all the alternatives and the worst alternative $\mathbf{x}_{W}$. Such comparisons can be collected into two sets as follows:
$A^{BO}=\left\{a_{Bi}\,|\,i\in R\right\}$ (12)
$A^{OW}=\left\{a_{iW}\,|\,i\in R\right\}$ (13)
Set $A^{BO}$ is called the ‘Best-to-Others’ set and its element $a_{Bi}$ shows the expert’s preference of the ‘best alternative’ $\mathbf{x}_{B}$ over alternative $\mathbf{x}_{i}$. Similarly, set $A^{OW}$ is called the ‘Others-to-Worst’ set and its generic element $a_{iW}$ represents the expert’s preference of alternative $\mathbf{x}_{i}$ over the worst alternative $\mathbf{x}_{W}$. The preferences can be expressed using a number between 1 to 9 (or other scales), i.e. $a_{Bi},a_{iW}\in\{1,\ldots,9\}$, where 1 represents ‘equality’ and 9 represents the ‘extreme preference’ of the former alternative over the latter. In a situation expressing full rationality $V(\mathbf{x}_{B})=a_{Bi}V(\mathbf{x}_{i})$ and $V(\mathbf{x}_{i})=a_{iW}V(\mathbf{x}_{W})$ for all $i\in R$. From this, it follows that
$\begin{cases}V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})=0\\ V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})=0\end{cases}\hskip 8.53581pt\forall i\in R$ (14)
As such, the objective would be to minimize the maximum deviation from the ideal situation (14), which can be formulated as the following linear optimization problem:
$\xi^{*}=$ minimize $\xi$ (BWD)
subject to $-\xi\leq V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})\leq\xi\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
$-\xi\leq V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})\leq\xi\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\in N\text{ and }\forall k$
$\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
$v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N\,.$
Solving (BWD), we can obtain the piecewise linear approximations of the attribute value functions and its optimum, $\xi^{*}$, is an indicator of compatibility of the preferences of the expert with the most suitable underlying value function $V_{L}$. If the results are deemed reliable, then the linear approximations can be used to rate the entire set of alternatives $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$. Coherently with the literature on preference disaggregation, we suggest that further analyses, especially sensitivity, be carried out to test the robustness of the results.
It must be pointed out that Barbati et al., (2024) already suggested a formulation compatible with (BWD)—see eq. (14) in their paper—with the only differences that (i) they employ the Deck-of-Cards method to rate alternatives (ii) they consider the final rating obtained from the Deck-of-Cards method instead of a set of comparisons $A^{BO}\cup A^{OW}$. Hereafter we will focus on (BWD) and propose methods to study consistency, compatibility and extreme rankings. Furthermore, we stress that we employed a min-max approach to remain coherent with the original BWM but, as shown also by Barbati et al., (2024), the optimization problem can be easily modified to accommodate other formulations, e.g., the sum of the errors.
3.1 A proposal to select the reference set
The choice of the reference set $R$ can be seen as a preprocessing phase, but this should not diminish its importance. While in existing disaggregation studies, the reference set is usually given, here, we consider some desiderata to propose a method to select the reference set:
  1. (D1):
    For all the attributes, the selected alternatives should cover all the segments into which the attribute ranges were divided. This ensures that there is preferential information on each segment of each attribute.
  2. (D2):
    Even considering the first desideratum, the reference set should be as informative as possible. Let us provide an example to justify our approach. If we consider two alternatives $a$ and $b$ such that $a$ outperforms $b$ with respect to all attributes, then knowing the degree of preference of $a$ over $b$ does not unveil information on the relative importance of attributes. Consider, instead, the two alternatives $c$ and $d$, and only the two attributes $j$ and $j^{\prime}$ such that $c$ scores better than $d$ with respect to $j$ and, vice versa, $d$ is better than $c$ with respect to $j^{\prime}$. In this case, knowing that $d$ is considered better than $c$ may give information on the greater importance of attribute $j^{\prime}$ over attribute $j$. Note also that $a$ Pareto dominates $b$, and that no Pareto dominance exists between $c$ and $d$. So, it may be wise to prefer a reference set where no contained alternative is dominated. Namely, we want to avoid stratified reference sets.
  3. (D3):
    The cardinality of the reference set should be minimized to decrease the cognitive load. This is coherent with current parsimonious approaches to the elicitation of preferences (Abastante et al., , 2019; Corrente et al., , 2024).
It is necessary to define auxiliary parameters and sets to formulate the optimization problem that puts together the three desiderata. We recall that each criteria is partitioned into $s_{j}$ (possibly equally wide) subintervals. For simplicity, we assume that $s_{j}$ is the same for each criterion and we call it $s$. At this point, one can simply construct a binary array $\mathbf{A}=(a_{ijk})_{m\times n\times s}\in\{0,1\}^{m\times n\times s}$ where
$a_{ijk}=\begin{cases}1,&\text{if the $i$th alternative covers the $k$th segment of the $j$th attribute}\\ 0,&\text{otherwise}\end{cases}$
We define $\mathcal{D}$ as the set of pairs $(i,i^{\prime})$ with $i<i^{\prime}$ and for which one alternative Pareto dominates another.
We propose a combinatorial optimization problem such that each segment is covered by at least $b\in\mathbb{N}_{+}$ alternatives and that in the selected set there is not any Pareto dominance relation between alternatives. That is, we can solve
minimize $\sum_{i\in M}y_{i}$ (15)
subject to $\sum_{i\in M}a_{i,j,k}y_{i}\geq b\hskip 14.22636pt\forall j,k$ (16)
$y_{i}+y_{i^{\prime}}\leq 1\hskip 14.22636pt\forall(i,i^{\prime})\in\mathcal{D}$ (17)
$y_{i}\in\{0,1\}\hskip 14.22636pt\forall i\in M$ (18)
[…]
It is also worth mentioning that the purpose of the proposed approach here is to find a non-dominated set as a reference point, not to find the Pareto optimal set of the alternative set. That is, the remaining alternatives which are not selected as the reference set are not necessarily dominated alternatives. Lastly, a cautionary note: it may be increasingly difficult to solve this problem for a growing number of splitting points. In fact, the greater number of segments to be covered conflicts with the constraint that the chosen reference alternatives cannot dominate each other.
3.2 Necessary relation from the BWD model
It could happen that the optimization problem (BWD) yields infinitely many optimal solutions. For this reason, we suggest an approach that tries to find the necessary ranking of the alternatives. That is, we propose an approach inspired by Robust Ordinal Regression (Corrente et al., , 2013). First, we call $\xi^{*}$ the optimal value of (BWD). For each pair of alternatives $(\mathbf{x}_{p},\mathbf{x}_{q})$, we maximize their value difference as follows:
$\delta_{pq}=$ maximize $V_{L}(\mathbf{x}_{p})-V_{L}(\mathbf{x}_{q})$ (19)
subject to $-\xi^{*}\leq V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
$-\xi^{*}\leq V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\in N\text{ and }\forall k$
$\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
$v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N$
Clearly, if $\delta_{pq}<0$, then for all optimal value functions, $V_{L}(\mathbf{x}_{p})-V_{L}(\mathbf{x}_{q})<0$. This implies that $\mathbf{x}_{q}$ is always valued higher than $\mathbf{x}_{p}$ across all optimal solutions. Hence, we can say that $\mathbf{x}_{q}$ is necessarily preferred to $\mathbf{x}_{p}$. In shorthand notation, we denote this necessary preference as $\mathbf{x}_{q}\succ_{\mathcal{N}}\mathbf{x}_{p}$. By repeatedly solving (19) for all pairs of distinct indices $({p},{q})\in M\times M$, we can elicit the necessary preference relation $\succ_{\mathcal{N}}$ across the entire set of alternatives. The results can be represented by a Hasse diagram.
3.3 Consistency analysis
In the previous section we recognized that the optimal value $\xi^{*}$ contains information on the discrepancy between the preferences expressed by the expert and the most suitable underlying model $V_{L}$. Hence, we can interpret $\xi^{*}$ as an index of compatibility. It is worth noting that compatibility should not be confused with the concept of (internal) consistency, which, in the case of the BWM, collapses into the condition $a_{Bj}a_{jW}=a_{BW}$ for all $j\in N$. Consider, for instance, that the preferences of the expert can be perfectly consistent, i.e. rational, and yet violate the principle of Pareto dominance between alternatives at the same time. This latter violation implies that the preferences cannot be fully compatible with the underlying model.
The conceptual difference between compatibility and consistency is shown in Figure 1. Consistency is an intrinsic property of the preferences expressed by an expert, while compatibility depends on both the preferences and the underlying model.
Figure 1: Compatibility and consistency.
Consequently, it might be more challenging to evaluate the significance of the value of the index $\xi^{*}$, especially since it is debatable whether this value should be minimized. In reality, $\xi^{*}$ is influenced by the design parameters, including the splitting points, of the model in question. An increase in the number of splitting points allows the underlying model to be more flexible and increases its alignment with preferences. However, artificially inflating the number of splitting points to improve compatibility could lead to overfitting issues.
We further note that, in the realm of preference disaggregation, the evaluation of internal consistency is a specific feature of the BWD, and possibly a unicum in the field of preference disaggregation. Other models, such as UTA and UTASTAR, require, as input, preference orders which are, by definition, transitive. Furthermore, they provide ordinal information and thus any analysis of cardinal information loses its grip. We conclude, also in light of Figure 1, that the analysis of compatibility is not enough and it should be coupled with an analysis of the consistency of the preferences. Starting from the principle that the level of inconsistency of preferences should not be model-dependent, therefore, we adopt the Ordinal Consistency Ratio ($OR$) and the Cardinal Consistency Ratio ($CR$) proposed by Liang et al., (2020) to quantify deviations from these consistency conditions in the BWD method.
The Ordinal Consistency Ratio focuses on the ranking of alternatives in pairwise comparisons, ensuring that the relative ordering of preferences is maintained, which is calculated as:
$OR=\max_{i}OR_{i}=\max_{i}\left\{\frac{1}{n}\sum_{k=1}^{n}F\left((a_{Bk}-a_{Bi})\cdot(a_{iW}-a_{kW})\right)\right\}$ (20)
where $F(x)$ is a step function, which is defined as:
$F(x)=\begin{cases}1,&\text{if }x<0,\\ 0.5,&\text{if }x=0\text{ and }(a_{Bk}-a_{Bi}\neq 0\text{ or }a_{iW}-a_{kW}\neq 0),\\ 0,&\text{otherwise.}\end{cases}$
The cardinal consistency check complements the ordinal consistency ratio by focusing on the magnitude of differences between pairwise comparisons. It examines whether the pairwise comparisons reflect the true scale of preferences, ensuring that the differences between alternatives are proportional to the decision maker’s intentions. To evaluate cardinal consistency, we use the input-based Cardinal Consistency Ratio ($CR$), which is defined as:
$CR=\begin{cases}\max_{i}\dfrac{|a_{Bi}a_{iW}-a_{BW}|}{|a_{BW}a_{BW}-a_{BW}|},&a_{BW}\geq 2\\ 0,&a_{BW}=0\end{cases}$ (21)
Identifying inconsistent judgments and understanding the extent of their deviation from full consistency is only part of the process. It is also crucial to determine acceptable thresholds for inconsistency for ensuring reliable decision outcomes in the BWD method. Since the consistency measurements in BWD mirror those in BWM, we adopt the thresholds for the input-based consistency ratio proposed by Liang et al., (2020). These thresholds act as benchmarks, where if the consistency ratios fall below these thresholds, the judgments are considered sufficiently consistent. However, if inconsistencies exceed these thresholds, an iterative process is employed to improve consistency. This process involves identifying the most inconsistent judgments, providing feedback to the decision-maker, revising the inconsistent comparisons, recalculating the $OR$ and $CR$, and repeating these steps until the judgments fall within the acceptable consistency range (Liang et al., , 2022). This approach ensures that the final preferences are refined and more consistent, leading to reliable and defensible decisions. Readers can refer to the original work (Liang et al., , 2022) for detailed steps and mathematical formulations.
3.4 Extreme ranking analysis
A number of sensitivity analysis tools already proposed for robust ordinal regression can be easily implemented for the BWD model. One example is the extreme ranking analysis Kadziński et al., (2012). In some decision-making contexts, it is often important to determine the maximum and minimum possible ranks that can be achieved by an alternative in the final ranking. This type of sensitivity analysis helps assess the stability of an alternative’s position. To determine these bounds, we solve two optimization problems for each alternative, always searching within the set of the optimal solutions of (BWD).
  • The outranking count ($\overline{d}_{i}$) represents the maximum number of alternatives that the given alternative $\mathbf{x}_{i}$ can outrank, at the same time. To determine this value, we solve the following optimization problem:
    $\overline{d}_{i}=$ maximize $\sum_{h\in M\setminus\{i\}}z_{h}$ (22)
    subject to $-\xi^{*}\leq V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
    $-\xi^{*}\leq V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
    $V_{L}(\mathbf{x}_{i})+\varepsilon\cdot z_{h}\geq V_{L}(\mathbf{x}_{h})-q_{h}\hskip 14.22636pt\forall h\in M\setminus\{i\}$
    $z_{h}+q_{h}=1\hskip 14.22636pt\forall h\in M$
    $v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\text{ and }\forall k$
    $\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
    $v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N$
    $z_{h},q_{h}\in\{0,1\}\hskip 14.22636pt\forall h\in M$
    where, again, $\varepsilon$ is an extremely small positive value. The optimization problem searches in the space of the optimal solutions of (BWD) the one that maximizes the number of alternatives that can simultaneously be judged worse that $\mathbf{x}_{i}$. This is done by maximizing the sum of the binary variables $z_{h}$, where $z_{h}=1$, if and only if, for that given compatible value function, $V_{L}(\mathbf{x}_{i})>V_{L}(\mathbf{x}_{h})$. The best possible rank ($\overline{r}_{i}$) of alternative $\mathbf{x}_{i}$ is then $\overline{r}_{i}=m-\overline{d}_{i}$, where $m$ represents the total number of alternatives being evaluated in the decision-making problem.
  • Similarly, the dominance count ($\underline{d}_{i}$) represents the maximum number of alternatives that can simultaneously be strictly preferred to $\mathbf{x}_{i}$. It is sufficient to solve the following problem to find $\underline{d}_{i}$:
    $\underline{d}_{i}=$ maximize $\sum_{h\in M\setminus\{i\}}z_{h}$ (23)
    subject to $-\xi^{*}\leq V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
    $-\xi^{*}\leq V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
    $V_{L}(\mathbf{x}_{h})+\varepsilon\cdot z_{h}\geq V_{L}(\mathbf{x}_{i})-q_{h}\hskip 14.22636pt\forall h\in M\setminus\{i\}$
    $z_{h}+q_{h}=1\hskip 14.22636pt\forall h\in M$
    $v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\text{ and }\forall k$
    $\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
    $v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N$
    $z_{h},q_{h}\in\{0,1\}\hskip 14.22636pt\forall h\in M$
    The worst possible rank ($\underline{r}_{i}$) of alternative $\mathbf{x}_{i}$ is then calculated as $\underline{r}_{i}=\underline{d}_{i}+1$.
By definition, $\overline{r}_{i}\leq\underline{r}_{i}$, and therefore we define the range of possible ranks as $[\overline{r}_{i},\underline{r}_{i}]$ for each alternative $i\in M$ with the goal of providing valuable insight into the stability and variability of the alternative’s ranking.
3.5 An interval-valued extension
Many decision analysis methodologies have been extended to accommodate uncertain and imprecise preference information. This need is particularly significant in the BWD method, where cardinal information is asked for holistic comparisons between alternatives, which are often more cognitively demanding than specific comparisons between criteria levels or in the form of trade-offs. Additionally, the interval-based approach allows experts to express uncertainty and imprecision more naturally by providing bounds on their judgments rather than a single precise value, and has the advantage that no assumption is made on the underlying probability distribution.
Using this approach, when an expert compares two alternatives $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, then we can assume that her judgement $\tilde{a}_{ij}=[{a}^{-}_{ij},{a}^{+}_{ij}]$ is an interval, meaning that, $V(\mathbf{x}_{i})/V(\mathbf{x}_{j})\in[{a}^{-}_{ij},{a}^{+}_{ij}]$. Equivalently,
$a_{ij}^{-}V(\mathbf{x}_{j})\leq V(\mathbf{x}_{i})\leq a_{ij}^{+}V(\mathbf{x}_{j})$ (24)
Nevertheless, it is possible that, due to conflicting preferences and inconsistencies, there may not exist an additive value function $V$ satisfying (24) and therefore one may want to relax its formulation using the slack variable $\xi$ as follows:
$-\xi+a_{ij}^{-}V(\mathbf{x}_{j})\leq V(\mathbf{x}_{i})\leq a_{ij}^{+}V(\mathbf{x}_{j})+\xi$ (25)
Consequently, we propose the following linear optimization problem as an extension of the BWD that can consider interval-valued comparisons.
$\xi_{I}^{*}=$ minimize $\xi$ (I-BWD)
subject to $-\xi+a_{Bi}^{-}V_{L}(\mathbf{x}_{i})\leq V_{L}(\mathbf{x}_{B})\leq a_{Bi}^{+}V_{L}(\mathbf{x}_{i})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{B\}$
$-\xi+a_{iW}^{-}V_{L}(\mathbf{x}_{W})\leq V_{L}(\mathbf{x}_{i})\leq a_{iW}^{+}V_{L}(\mathbf{x}_{W})+\xi\hskip 14.22636pt\forall{i}\in R\setminus\{W\}$
$v_{j}^{L}(x_{j}^{k})\leq v_{j}^{L}(x_{j}^{k+1})\hskip 14.22636pt\forall j\in N\text{ and }\forall k$
$\sum_{j=1}^{n}v_{j}(\overline{x}_{j})=1$
$v_{j}(\underline{x}_{j})=0\hskip 14.22636pt\forall j\in N$
$\xi\geq 0.$
Note that (I-BWD) is a generalization of (BWD) and when the preference information is given in the form of real numbers, the former collapses into the latter. Indeed, all the steps proposed to test the robustness of the real-valued BWD method—i.e. optimization problems (19), (22), and (23)—can be extended to the interval case by replacing the constraints
$\begin{cases}-\xi^{*}\leq V_{L}(\mathbf{x}_{B})-a_{Bi}V_{L}(\mathbf{x}_{i})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{B\}\\ -\xi^{*}\leq V_{L}(\mathbf{x}_{i})-a_{iW}V_{L}(\mathbf{x}_{W})\leq\xi^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{W\}\end{cases}$ (26)
with the generalized interval constraints:
$\begin{cases}-\xi_{I}^{*}+a_{Bi}^{-}V_{L}(\mathbf{x}_{j})\leq V_{L}(\mathbf{x}_{i})\leq a_{Bi}^{+}V_{L}(\mathbf{x}_{j})+\xi_{I}^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{B\}\\ -\xi_{I}^{*}+a_{iW}^{-}V_{L}(\mathbf{x}_{j})\leq V_{L}(\mathbf{x}_{i})\leq a_{iW}^{+}V_{L}(\mathbf{x}_{j})+\xi_{I}^{*}\hskip 14.22636pt\forall{i}\in R\setminus\{W\}\end{cases}$ (27)
Indeed, BWD can be seen as a restriction of I-BWD. That is, given a problem, if all the real-valued preferences lie within the interval-valued ones, then $\xi^{*}_{I}\leq\xi^{*}$.
Proposition 1.
If $a_{Bi}\in[a_{Bi}^{-},a_{Bi}^{+}]$, $a_{iW}\in[a_{iW}^{-},a_{iW}^{+}]$ and the consequence vectors are the same for both models, then, $\xi^{*}_{I}\leq\xi^{*}$.
Proof.
Let us consider the case of the comparisons between the best and the others. In the I-BWD, the first family of constraints is a linearization of
$\xi_{I}\geq\max\left\{V(x_{B})-a^{+}_{Bi}V(x_{i}),a^{-}_{Bi}V(x_{i})-V(x_{B}),0\right\}$ (28)
In the BWD, if we consider the real value $a_{Bi}$ as a degenerate interval, i.e. $a_{Bi}\in[a_{Bi},a_{Bi}]$, then we can write
$\xi\geq\max\left\{V(x_{B})-a_{Bi}V(x_{i}),a_{Bi}V(x_{i})-V(x_{B}),0\right\}.$ (29)
As we know that $a_{Bi}^{-}\leq a_{Bi}\leq a_{Bi}^{+}$ and $V(\mathbf{x}_{i})\geq 0$ for all $i\in R$ we deduce
$V(x_{B})-a^{+}_{Bi}V(x_{i})$ $\leq V(x_{B})-a_{Bi}V(x_{i})$
$a^{-}_{Bi}V(x_{i})-V(x_{B})$ $\leq a_{Bi}V(x_{i})-V(x_{B})$
and therefore the right-hand side of (29) cannot be smaller than the right-hand side of (28). We skip the formal proof, but a similar line of reasoning can be applied to the second family of constraints for both (BWD) and (I-BWD). If we consider that $\xi\geq 0$ naturally holds in (BWD), then we can see that all other constraints are the same, guaranteeing the same freedom in the definition of $V$. Hence, due to the minimization of both $\xi_{I}$ and $\xi$, we obtain $\xi_{I}^{*}\leq\xi^{*}$. ∎
4 A case study in logistics performance evaluation
We tested the BWD on the data collected and used by Rezaei et al., (2018) to rank countries according to a composite indicator of logistic performance. In their study, data from the World Bank was used to assess the performance of the World countries with respect to the following criteria: Customs; Infrastructure; Services; Timeliness; Tracking and tracing; International shipment. For the sake of simplicity and ease of interpretation of the results, we only focused on European countries.
We used the minimum and maximum values of each attribute as lower and upper levels of each attribute. That is, given the decision matrix $\mathbf{X}=(x_{ij})_{m\times n}$ and considering that all six criteria were originally expressed as benefit criteria, we defined
$\underline{x}_{j}=\min_{i\in M}x_{ij}\;\;\text{ and }\;\;\overline{x}_{j}=\max_{i\in M}x_{ij}$
Reference set selection and preference elicitation
One splitting point was allowed for each criterion set and was positioned midway in the interval $[\underline{x}_{j},\overline{x}_{j}]$. Next, we solved the set covering-inspired optimization problem (15)–(18) to find a mutually non-dominated minimum covering set. This set contained only three countries and was taken as a good initial suggestion. Given the possibility of asking for more comparisons, we added two more alternatives to better explore the lower part of the ranking. The final reference set $R$ was defined as
[…]
By adding the coverage provided by each alternative for each criteria segment, we arrive at the following matrix.
$\mathbf{A}^{R}=\left(\sum_{i\in R}a_{ijk}\right)_{j\in N,k\in\{1,2\}}\\ =\underbrace{\begin{pmatrix}0&1\\ 1&0\\ 1&0\\ 1&0\\ 1&0\\ 0&1\end{pmatrix}+\begin{pmatrix}1&0\\ 0&1\\ 0&1\\ 0&1\\ 0&1\\ 0&1\end{pmatrix}+\begin{pmatrix}0&1\\ 0&1\\ 0&1\\ 1&0\\ 0&1\\ 1&0\end{pmatrix}}_{=\begin{pmatrix}1&2\\ 1&2\\ 1&2\\ 2&1\\ 1&2\\ 1&2\end{pmatrix}}+\begin{pmatrix}1&0\\ 0&1\\ 1&0\\ 1&0\\ 0&1\\ 0&1\end{pmatrix}+\begin{pmatrix}1&0\\ 1&0\\ 1&0\\ 1&0\\ 1&0\\ 1&0\end{pmatrix}=\begin{pmatrix}3&2\\ 2&3\\ 3&2\\ 4&1\\ 2&3\\ 2&3\end{pmatrix}$
As each entry of the matrix $\mathbf{A}^{R}$ is at least one, then we know that the alternatives indexed in $R$ cover all attribute levels. We can see that, as expected, the first three matrices are sufficient to cover all criteria levels and that none of the first three alternatives dominate the other. This corroborates the validity of the optimization problem (15)–(18). Next, an expert used the 1–9 scale and provided preference information on the reference alternatives by means of the $A^{BO}$ and $A^{OW}$ sets presented in Table 1.
Estonia Hungary Latvia Greece Moldova
$A^{BO}$ ($B$=Estonia) 1 3 4 5 8
$A^{OW}$ ($W$=Moldova) 8 5 3 4 1
Table 1: Best-to-Others ($A^{BO}$) and Others-to-Worst ($A^{OW}$) pairwise comparison vectors.
Once we acquired the preferences in $A^{BO}$ and $A^{OW}$, we verified their consistency to detect possibly inaccurate and unreliable preference information. This involves checking both ordinal and cardinal consistency, and making adjustments if necessary.
Consistency analysis
First, we assess the cardinal consistency by calculating the Input-based Consistency Ratio ($CR$) according to formula (21). In this case, $CR=0.214$, which is less than the threshold (the input-based consistency threshold table can be checked in (Liang et al., , 2020)) of 0.284 for 5 alternatives and $a_{BW}=8$. This indicates that the cardinal inconsistency of the provided preferences is tolerable, allowing us to proceed to the next step.
Using the Ordinal Consistency Ratio ($OR$) as formulated in eq. (20), we found $OR=0.2$, indicating a violation of ordinal consistency. The local Ordinal Consistency Ratios ($OR_{i}$), shown in Table 2, help locate specific inconsistencies. In this case, the preferences provided by the expert on Latvia are not ordinally consistent with those expressed on Greece.
Estonia Hungary Latvia Greece Moldova $OR_{i}$
Estonia 0 0 0 0 0 0
Hungary 0 0 0 0 0 0
Latvia 0 0 0 1 0 0.2
Greece 0 0 1 0 0 0.2
Moldova 0 0 0 0 0 0
Table 2: The ordinal consistency check table: “1” indicates an ordinal inconsistency between the corresponding alternatives, while “0” indicates consistency.
After identifying and locating inconsistencies in the preferences, we re-engaged with the expert to review and reconsider her judgments. Upon reflection, the expert acknowledged the inconsistencies and proceeded to modify the preferences using the consistency improvement process from the work of Liang et al., (2022). Using the same approach, we derived the admissible ranges for $A^{BO}$ and $A^{OW}$, as shown in Table 3, and the visualization of the improving ranges, shown in Figure 2.
Alternatives Estonia Hungary Latvia Greece Moldova
Original [1, 8] [1, 8] [1, 8] [1, 8] [1, 8]
Acceptable ranges $A^{BO}$ [1, 2.99] [1, 4.78] [1, 7.97] [1, 5.98] [1, 8]
$A^{OW}$ [1, 8] [1, 7.97] [1, 5.98] [1, 4.78] [1, 2.99]
Improving ranges $A^{BO}$ [1, 1] [1.6, 3] [2.67, 4] [2, 5] [8, 8]
$A^{OW}$ [8, 8] [3.6, 5] [1.67, 3] [1, 4] [1, 1]
Table 3: The admissible ranges for revision.
Figure 2: Improving ranges, in green, for $A^{BO}$ and $A^{OW}$.
With the help of the visualization of the ranges for improvement, the expert can easily adjust the inconsistent comparison. After consideration, the expert made some minor revisions on the preferences. The modified sets $A^{BO}$ and $A^{OW}$ are reported in Table 4.
Estonia Hungary Latvia Greece Moldova
$A^{BO}$ ($B$=Estonia) 1 3 4 4 8
$A^{OW}$ ($W$=Moldova) 8 5 3 3 1
Table 4: Revised Best-to-Others ($A^{BO}$) and Others-to-Worst ($A^{OW}$) pairwise comparison vectors.
After revising the preferences, we rechecked their consistency using both the ordinal and cardinal consistency indices. The revised preferences showed an Ordinal Consistency Ratio $OR=0$, indicating full ordinal consistency. The Cardinal Consistency Ratio is $CR=0.125$, which is below the threshold of $0.284$ for 5 alternatives and $a_{BW}=8$. This confirms that the judgments are now acceptably consistent and represent an improvement over the initial preferences before revision.
Optimization results and compatibility
Preferences were used as inputs in the BWD optimization problem, and an optimal value $\xi^{*}=0.030689$ was found. This means that even the best fitting model $V_{L}$ cannot comply exactly with the expert’s judgments. This should be allowed, since the cardinal preferences of an expert expressed as real numbers can hardly ever fully reflect an underlying piecewise linear model.
The full analysis revealed that $\succ_{\mathcal{N}}$ is almost a total order relation. This is confirmed in Figure 3, showing very low variability in the set of rankings obtained through optimization problems (22) and (23). Only one swap is possible, between Greece and Slovenia. The results shown in Figure 3 seem to hint at a good precision of the outcome and further tests that we conducted, but we avoided reporting, showed any level of incompatibility, i.e. $\xi^{*}>0$, very often leads to an extremely stable ranking.
Figure 3: Rankings of European countries based on BWD (on the left) and I-BWD (on the right).
Interval-BWD
We also tested the I-BWD on the same case study, together with the analysis of dominance relations and maximum and minimum ranks. In this case, the expert information is summarized in Table 5. Note that the improving ranges shown in Table 3 can be used to guide experts to extend their original real-valued preferences to interval ones in a way that may decrease their inconsistency.
Estonia Hungary Latvia Greece Moldova
$\tilde{A}^{BO}$ ($B$=Estonia) $[1,1]$ $[2,3]$ $[3,4]$ $[4,5]$ $[7,9]$
$\tilde{A}^{OW}$ ($W$=Moldova) $[7,9]$ $[4,5]$ $[2,4]$ $[3,4]$ $[1,1]$
Table 5: The two sets of interval-valued comparisons: $\tilde{A}^{BO}$ and $\tilde{A}^{OW}$.
In the case of I-BWD, the imprecise information provided by the expert was sufficient to achieve full compatibility between the preferences and the underlying model, i.e. $\xi^{*}_{I}=0$. Consequently, the necessary relation $\succ_{\mathcal{N}}$ is less informative, as shown in the Hasse diagram in Figure 4.
Figure 4: Necessary ranking $\succ_{\mathcal{N}}$ of European countries obtained with the interval-valued BWD. Countries are labeled using the ISO 3166-1 alpha-2 standard.
Sensitivity analysis
Results of the extreme ranking analysis are presented in the right-hand side of Figure 3 and confirms the previous intuition on the much greater variability of the attainable ranks. To quantify the imprecision due to the existence of multiple compatible rankings, we introduce the following heuristic index
$U=\frac{1}{m}\sum_{i\in M}\frac{\underline{r}_{i}-\overline{r}_{i}}{m-1}\in[0,1]$
where each term of the sum is the “spread” of the possible ranks that the $i$th alternative can attain divided by the maximum possible range, going from $1$, the best rank, to $m$, the lowest possible attainable rank. The scaling factor preceding the sum normalizes the quantity into the interval $[0,1]$. In the case of the I-BWD, we have $U=0.05938$, whereas, for the real-valued BWD it was $U=0.00135$.
Computational feasibility
The analysis for the I-BWD, including the necessary relation $\succ_{N}$ represented in Figure 4 and the maximum and minimum ranks in Figure 3 was carried out with Wolfram Mathematica 14.0.0, with a total computational time of 16.65 seconds on an Intel(R) Core(TM) i7-8550U CPU @ 1.80GHz with 16Gb RAM. Optimization problems were solved using Gurobi 11.0.1. If the problem was run for the entire set considered by Rezaei et al., (2018)—165 World countries—then the computational time would still remain under 3 minutes. These results seem to support the computational feasibility of the proposed approach. Furthermore, out of the 16.65 seconds employed for the case of European countries, 13.53 seconds were spent in solving problem (19), which, as it shall be solved for all distinct pairs of indices $(p,q)$, has quadratic complexity with respect to $m$. On the contrary, the optimization problems used to find the best and worst attainable ranks have linear complexity with respect to $m$. Hence, if the optimization problems used to find $\succ_{\mathcal{N}}$ were omitted, then the entire procedure would have a much greater scalability.
5 Conclusions
The Best-Worst Disaggregation Method (BWD) offers an advancement in the field of Multi-Criteria Decision Analysis (MCDA) by combining the principles of the Best-Worst Method (BWM) with the disaggregation approach. This method addresses the challenges inherent in traditional aggregation and disaggregation techniques by leveraging the strengths of both to create a robust and reliable decision-making framework. It is a valuable tool in scenarios where experts have a clear understanding of alternatives but may not have predefined value functions.
BWD’s core advantage lies in its ability to derive attribute value functions directly from the experts’ two pairwise comparison vectors of alternatives. This process ensures that the resulting weights and value functions align closely with the actual preferences of the experts, leading to more reliable and meaningful outcomes. BWD reduces the cognitive load on experts and improves the reliability of derived models by simplifying the comparison process and ensuring consistency through admissible ranges for iterative visual refinements. Moreover, BWD’s structured approach to preference elicitation and consistency improvement ensures that the decision-making process remains transparent.
Besides, we proposed a method to select the reference set that ensures comprehensive coverage of attribute levels, avoids Pareto dominance among alternatives, and minimizes cognitive load on the decision-makers. This approach enhances the representativeness and effectiveness of the reference set, contributing to more accurate and reliable preference modeling.
Furthermore, we extended the BWD method to handle interval-valued preferences, allowing for the accommodation of uncertainty or imprecise information in expert judgments. This extension broadens the applicability of BWD in real-world decision-making scenarios where exact numerical assessments may not be feasible.
Future research on BWD could adapt it to accommodate group decision-making scenarios by developing techniques to aggregate individual preferences and address potential conflicts; adapting BWD for use in dynamic environments where decision criteria and alternatives may change over time; enhancing the computational efficiency of BWD to handle larger sets of criteria and alternatives by exploring advanced optimization techniques and parallel computing strategies; investigating methods to minimize biases in expert judgments, such as incorporating techniques from behavioral decision theory or utilizing machine learning or artificial intelligence to identify and correct inconsistencies; conducting practical case studies in various domains to validate the effectiveness of BWD and identify specific areas for improvement. Finally, in this manuscript we employed the well-known 1–9 scale, but other methods to collect statements on alternatives, such as those used in GRIP (Figueira et al., , 2009), may be incorporated.
Acknowledgments
We thank Majid Mohammadi for his comments on early versions of this manuscript.